y'(0)=1
La placE TRANSFORMto solve y - 2ylty = f(t), yco)=7, yie f(t) = soi tl5 It-5, t25 SHOW all STEPS
Answer #1

Solution y - 2 y + y = f(t) t 5 f(t)= 2 • , #25 t-5, +7,5 Taking Laplace transform, we get L{ y – 2y ty} = L {f His = f est